OpenStudy (anonymous):

find all complex numbers z satisfying:

OpenStudy (anonymous):

\[(2z+1+\sqrt{3}/2z+1-\sqrt{3})^{3}+(2z+1+\sqrt{3}/2z+1-\sqrt{3})^{2}+ (2z+1+\sqrt{3}/2z+1-\sqrt{3})\]

OpenStudy (anonymous):

\[(2z+1+\sqrt{3}/2z+1-\sqrt{3})\] < part that was cut off

OpenStudy (mr.math):

Satisfying what exactly? That's just an expression!

OpenStudy (mr.math):

Does all that equal to 0, for example?

OpenStudy (anonymous):

sorry! there should be a +1 added to the end as well, and equals 0. apologies!

OpenStudy (mr.math):

So the equation is \[\small (\frac{2z+1+\sqrt{3}}{2z+1-\sqrt{3}})^3+(\frac{2z+1+\sqrt{3}}{2z+1-\sqrt{3}})^2+(\frac{2z+1+\sqrt{3}}{2z+1-\sqrt{3}})+a+1=0?\]

OpenStudy (anonymous):

yes thats it

OpenStudy (anonymous):

i wasn't able to write it all out for some reason it messed up!

OpenStudy (anonymous):

without the a, thats not involved... that was just me saying there is +1 at the end